A triangle whose height is 11 less than its base has an area of 13. What are the measures of the base and height?
Answer by richwmiller(17219)

h=b-11
1/2b*h=13
b*h=26
b*(b-11)=26
b^2-11b-26=0
(b-13)*(b+2)=0
b=13
h=2
check
1/2*b*h=13
1/2*2*13=13
13=13
ok